# On the geometry of spaces of filtrations on local rings

### v1

@@ Abstract.
We study the geometry of spaces of fitrations on a Noetherian local domain. We introduce a metric $d_{1}$ on the space of saturated filtrations, inspired by the Darvas metric in complex geometry, such that it is a geodesic metric space. In the toric case, using Newton-Okounkov bodies, we identify the space of saturated monomial filtrations with a subspace of $L^{1}_{\mathrm{loc}}$. We also consider several other topologies on such spaces and study the semi-continuity of the log canonical threshold function in the spirit of Kollár-Demailly. Moreover, there is a natural lattice structure on the space of saturated filtrations, which is a generalization of the classical result that the ideals of a ring form a lattice.
@@ 1. Introduction
As a generalization of ideals, graded sequences of ideals have been extensively studied in the past few years, and have played an important roles in many areas of algebraic geometry, see for example [ELS03, JM12, Cut13] and [BFJ08, BdFF12, BFJ14, Li18, Blu18, Liu18, Xu20, XZ21, LXZ22].
Filtrations on a local ring $(R,\mathfrak{m})$ are the continuously indexed version of graded sequences of ideals. To be more precise, an $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ on $R$ is a collection $\{\mathfrak{a}_{\lambda}\}_{\lambda\in\mathbb{R}_{>0}}$ of $\mathfrak{m}$-primary ideals, which is decreasing, multiplicative and left continuous. The three conditions mean $\mathfrak{a}_{\lambda}\subset\mathfrak{a}_{\mu}$ for $\lambda>\mu$, $\mathfrak{a}_{\lambda}\cdot\mathfrak{a}_{\mu}\subset\mathfrak{a}_{\lambda+\mu}$ and $\mathfrak{a}_{\lambda-\epsilon}=\mathfrak{a}_{\lambda}$ for $0<\epsilon\ll 1$, respectively.
In this paper, we study the geometry of certain spaces of such filtrations and prove some structural results. The first main result is that there is a pseudometric $d_{1}$ on the set of $\mathfrak{m}$-filtrations with positive multiplicity, which is an analogue of the Darvas metric introduced in [Dar15] in complex geometry (see [BJ21] for the non-archimedean setting), such that a suitable retraction, the set of saturated filtrations, is a geodesic metric space. In the toric setting, the subspace of saturated monomial filtrations can be identified with the space of cobounded convex sets of the dual cone with the symmetric difference metric, which is naturally a subspace of $L^{1}_{\mathrm{loc}}$. We also introduce another metric, denoted by $d_{\infty}$, and several weak topologies, and study the continuity properties of the log canonical threshold function, following Demailly-Kollár [DK01]. Several spaces of filtrations have a natural structure of a lattice, generalizing a classical result that the ideals of a ring form a modular lattice.
@@ 1.1.1. The Darvas metric, multiplicities and saturations
Given a compact Kähler manifold $(X,\omega)$, Darvas [Dar15] introduced various Finsler metrics on the space $\mathcal{H}$ of smooth Kähler potentials, generalizing the metric studied in [Mab87]. One of the metrics, known as the Darvas metric $d_{1}$, has found numerous applications in complex geometry. In particular, the geodesic metric space $(\mathcal{H},d_{1})$ and its completion, $(\mathcal{E}^{1},d_{1})$, the space of Kähler potentials of finite energy, are closely related to Kähler-Einstein metrics. See for example [DR17, BBEGZ19, DL20]. A comprehensive survey on this metric space and its applications, particularly in relation to Tian’s Properness Conjecture, can be found in [Dar19].
More recently, Boucksom and Jonsson [BJ21] developed a similar construction in the non-archimedean setting , where they introduced a pseudo-metric $d_{1}$, also referred to as the Darvas metric, on the space $\mathcal{N}$ of norms on the section ring $R(X,-mrK_{X})$ of a projective Fano variety $X$. Geodesics on this space have been independently studied by Reboulet [Reb22] and by Blum, Liu, Xu and Zhuang [BLXZ21], and these studies have found successful applications in the algebraic theory of $K$-stability and Kähler-Ricci soliton degenerations. For some related constructions, see also [Reb20, Wu22, Fin23].
@@
$\mathrm{vol}(\mathfrak{a}_{\bullet})\coloneqq\lim_{\mathbb{Z}_{>0}\ni m\to\infty}\frac{\ell(R/\mathfrak{a}_{m})}{m^{n}/n!}=\lim_{\mathbb{Z}_{>0}\ni m\to\infty}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}},$
@@
$d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\coloneqq 2\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}),$ (1.1)
where $\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}$ is defined termwise. (1.1) is a similar expression to the Darvas metrics in the global case [Dar15, BJ21].
Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}_{R,\mathfrak{m}}$, the geodesic $\mathfrak{a}_{\bullet,t}$ between them, introduced in [XZ21, BLQ24], is a segment of $\mathfrak{m}$-filtrations for $t\in[0,1]$ defined by
$\mathfrak{a}_{\lambda,t}\coloneqq\sum_{\mu,\nu\geq 0,\ (1-t)\mu+t\nu=\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}.$ (1.2)
The definition is a local analogue of the geodesics studied in [BLXZ21].
Our first main result is that the constructions above endow the Hausdorff quotient of $(\mathrm{Fil}_{R,\mathfrak{m}},d_{1})$ with the structure of a geodesic metric space.
@@
Let $(R,\mathfrak{m})$ be a Noetherian local domain that is analytically irreducible. Denote the set of all linearly bounded $\mathfrak{m}$-filtrations on $R$ by $\mathrm{Fil}_{R,\mathfrak{m}}$. Then
@@
    the function $d_{1}$ given by (1.1) is a pesudometric on $\mathrm{Fil}_{R,\mathfrak{m}}$,
@@
Here $\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ is the subset of $\mathrm{Fil}_{R,\mathfrak{m}}$ consisting of saturated $\mathfrak{m}$-filtrations. For the definition of saturated filtrations, see section 2.2. For definitions related to metric spaces, see section 2.6.
@@ 1.1.2. The toric case
Let $\mathbbm{k}$ be a field, $N$ a free abelian group of rank $n\geq 1$ and $M=N^{*}$ its dual. Let $\sigma\subset N_{\mathbb{R}}=N\otimes_{\mathbb{Z}}\mathbb{R}$ be a strongly convex rational polyhedral cone of maximal dimension. This gives us an affine normal toric variety $X_{\sigma}=\mathrm{Spec}R_{\sigma}=\mathrm{Spec}\mathbbm{k}[\sigma^{\vee}\cap M]$ with a unique torus invariant point $x$, where $\sigma^{\vee}\subset M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$ is the dual cone of $\sigma$. Let $R$ be the local ring of $X$ at $x$ and $\mathfrak{m}$ its maximal ideal.
Let $\mathrm{Fil}^{s\mathrm{mon}}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ be the subset of all monomial filtrations (i.e. filtrations $\mathfrak{a}_{\bullet}$ such that every $\mathfrak{a}_{\lambda}$ is monomial). Via the Okounkov body construction, we identify the metric space $(\mathrm{Fil}^{S,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ with a subspace of the space of co-bounded subsets of $\sigma^{\vee}$ with the symmetric difference metric (sometimes also called the Fréchet-Nykodym-Aronszyan metric).
@@
    the metric space $(\mathrm{Fil}^{s,\mathrm{mon}}_{X,x},d_{1})$ is complete.
In the above theorem, $\mathcal{P}(\sigma^{\vee})$ denotes the set of all closed convex subsets of $\sigma^{\vee}$ with bounded complement, and the metric $d$ is defined to be the Euclidean volume of the symmetric difference of two subsets. For more details, we refer to Section 3.4. Note that $P\mapsto\chi_{P}$ gives an embedding $\mathcal{P}(\sigma^{\vee})\to L^{1}_{\mathrm{loc}}(\sigma^{\vee})$.
@@ 1.1.3. The supnorm metric and homogeneous norms
As in the global case [BJ21], there is also a correspondence between filtrations and norms in the local setting, see Definition 2.7 for the definition of norms and Definition-Lemma 2.8, Lemma 2.12 and Lemma 2.29 for the dictionary. Despite the equivalence, it is sometimes more convenient to use one of the languages.
As an example, we now consider another pseudometric $d_{\infty}$ on the space $\mathcal{N}_{R,\mathfrak{m}}$ of $\mathfrak{m}$-norms on $R$, which behaves like the uniform metric on the space of continuous functions on a compact space. This can also be viewed as the local analogue of the $d_{\infty}$ metric introduced in [BJ21].
@@
$d_{\infty,\rho}(\chi,\chi^{\prime})\coloneqq\limsup_{\lambda\to\infty}\sup_{\rho(f)\geq\lambda}\frac{|\chi(f)-\chi^{\prime}(f)|}{\rho(f)}.$ (1.3)
@@
1.2. Weak topologies and semicontinuity of log canonical thresholds
@@
Another point of view is to consider a filtration $\mathfrak{a}_{\bullet}$ as a function on certain valuation spaces by $v\mapsto v(\mathfrak{a}_{\bullet})$, where $v(\mathfrak{a}_{\bullet})\coloneqq\lim_{\lambda\to\infty}\frac{v(\mathfrak{a}_{\lambda})}{\lambda}$. We call the topology defined this way using the space $\mathrm{Val}^{+}_{R,\mathfrak{m}}$ of valuations with positive volume the $+$-topology.
We prove that the log canonical threshold satisfies certain semicontinuity on the space of filtrations, and that it is locally Lipschitz continuous with respect to the $d_{\infty}$-topology, which is similar to [DK01, Theorem 3.3].
@@
Let $(X,x)=(\mathrm{Spec}R,\mathfrak{m})$ be a klt singularity over a field $\mathbbm{k}$ of characteristic 00. Then
@@
    if $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}$ converges to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$ in the $+$-topology, then
@@ 1.3. Miscellaneous results about filtrations
The spaces of filtrations are equipped with a natural partial order by inclusion, that is, $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$ if and only if $\mathfrak{a}_{\lambda}\subset\mathfrak{b}_{\lambda}$ for any $\lambda\in\mathbb{R}_{>0}$. One can similarly define the intersection of two filtration termwisely, and in Definition-Lemma 2.3 (Definition-Lemma 2.22, resp.), we define the join $\vee$ (saturated join $\vee_{s}$, resp.), which is the analogue of the sum of two ideals. The next theorem asserts that these operations define natural lattice structures on the spaces of $\mathfrak{m}$-filtrations.
@@
This theorem, together with Theorem 1.1, can be viewed as a higher-dimensional analogue of the results of [FJ04].
We will frequently use the following characterization for saturated filtrations, which can be viewed as an analogue of the Berkovich Maximum Modulus Principle. See also Lemma 2.32.
@@
As an example of the applications, when $\mathfrak{a}$ is an $\mathfrak{m}$-primary ideal, we prove that the saturation of the $\mathfrak{a}$-adic filtration is determined by the set of Rees valuations of $\mathfrak{a}$, as expected.
@@
Organization of the paper. In Section 2, we recall some definitions and basic properties that will be needed. Along the way we also prove some basic facts about filtrations. In Section 3, we introduce the Darvas metric $d_{1}$ and consider its properties. In Section 4, we define the metric(s) $d_{\infty}$, and compare the different topologies on the space of filtrations. In Section 5 we prove the main results Theorem 1.1 through Theorem 1.5. In Section 6, we discuss the relation of the results in this paper with results in the global case, give several further examples, and propose some related questions.
@@
Acknowledgement. A large portion of this paper is derived from the author’s doctoral thesis. The author would like to thank his advisor, Professor Chenyang Xu, for his constant support, encouragement and numerous inspiring conversations. Part of this work is inspired by an earlier collaboration with Harold Blum and Yuchen Liu. The author owes them special thanks for their insights on the topic.
The author would like to thank Zhiyuan Chen, Colin Fan, Yujie Luo, Tomasso de Fernex, Rémi Reboulet, Xiaowei Wang and Ziquan Zhuang for fruitful discussions. The author would also like to thank Tamás Darvas, Jingjun Han, Jihao Liu, Minghao Miao, Junyao Peng, Linsheng Wang, Lingyao Xie, Qingyuan Xue, Tong Zhang and Junyan Zhao for kind comments. Special thanks are also due to Rémi Reboulet and Mattias Jonsson for reviewing a preliminary version of this paper and offering valuable suggestions. Part of this work is done while the author visited the Northwestern University, the University of Utah, Fudan University and BICMR at Peking University. The author would like to thank their hospitality and the amazing environment they provided.

### v2

@@ Abstract.
We study the geometry of spaces of filtrations on a Noetherian local domain. We introduce a metric $d_{1}$ on the space of saturated filtrations, inspired by the Darvas metric in complex geometry, such that it is a geodesic metric space. In the toric case, using Newton-Okounkov bodies, we identify the space of saturated monomial filtrations with a subspace of $L^{1}_{\mathrm{loc}}$. We also consider several other topologies on such spaces and study the semi-continuity of the log canonical threshold function in the spirit of Demailly-Kollár. Moreover, there is a natural lattice structure on the space of saturated filtrations, which is a generalization of the classical result that the ideals of a ring form a lattice.
@@ 1. Introduction
As a generalization of ideals, graded sequences of ideals have been extensively studied in the past few years, and have played an important role in many areas of algebraic geometry; see for example [ELS03, JM12, Cut13] and [BFJ08, BdFF12, BFJ14, Li18, Blu18, Liu18, Xu20, XZ21, LXZ22].
Filtrations on a local ring $(R,\mathfrak{m})$ are the continuously indexed version of graded sequences of ideals. To be more precise, an $\mathfrak{m}$-filtration $\mathfrak{a}_{\bullet}$ on $R$ is a collection $\{\mathfrak{a}_{\lambda}\}_{\lambda\in\mathbb{R}_{>0}}$ of $\mathfrak{m}$-primary ideals, which is decreasing, multiplicative, and left continuous. The three conditions mean $\mathfrak{a}_{\lambda}\subset\mathfrak{a}_{\mu}$ for $\lambda>\mu$, $\mathfrak{a}_{\lambda}\cdot\mathfrak{a}_{\mu}\subset\mathfrak{a}_{\lambda+\mu}$ and $\mathfrak{a}_{\lambda-\epsilon}=\mathfrak{a}_{\lambda}$ for $0<\epsilon\ll 1$, respectively.
In this paper, we study the geometry of certain spaces of such filtrations and prove some structural results. The first main result is that there is a pseudometric $d_{1}$ on the set of $\mathfrak{m}$-filtrations with positive multiplicity, which is an analogue of the Darvas metric introduced in [Dar15] in complex geometry (see [BJ25] for the non-archimedean setting), such that under a suitable retraction, the set of saturated filtrations is a geodesic metric space. In the toric setting, the subspace of saturated monomial filtrations can be identified with the space of cobounded convex sets of the dual cone with the symmetric difference metric, which is naturally a subspace of $L^{1}_{\mathrm{loc}}$. We also introduce another metric, denoted by $d_{\infty}$, and several weak topologies, and study the continuity properties of the log canonical threshold function, following Demailly-Kollár [DK01]. Several spaces of filtrations have a natural structure of a lattice, generalizing a classical result that the ideals of a ring form a modular lattice.
@@ 1.1.1. The Darvas metric, multiplicities and saturations
Given a compact Kähler manifold $(X,\omega)$, Darvas [Dar15] introduced various Finsler metrics on the space $\mathcal{H}$ of smooth Kähler potentials, generalizing the metric studied in [Mab87]. One of the metrics, known as the Darvas metric $d_{1}$, has found numerous applications in complex geometry. In particular, the geodesic metric space $(\mathcal{H},d_{1})$ and its completion, $(\mathcal{E}^{1},d_{1})$, the space of Kähler potentials of finite energy, are closely related to Kähler-Einstein metrics. See, for example, [DR17, BBEGZ19, DL20]. A comprehensive survey on this metric space and its applications, particularly in relation to Tian’s Properness Conjecture, can be found in [Dar19].
More recently, Boucksom and Jonsson [BJ25] developed a similar construction in the non-archimedean setting, where they introduced a pseudometric $d_{1}$, also referred to as the Darvas metric, on the space $\mathcal{N}$ of norms on the section ring $R(X,L)$ of a polarized variety $(X,L)$. Geodesics on this space have been independently studied by Reboulet [Reb22] and by Blum, Liu, Xu and Zhuang [BLXZ21], and these studies have found successful applications in the algebraic theory of $K$-stability and Kähler-Ricci soliton degenerations. For some related constructions, see also [Reb20, Wu22, Fin25a].
@@
$\mathrm{e}(\mathfrak{a}_{\bullet})\coloneqq\lim_{\mathbb{Z}_{>0}\ni m\to\infty}\frac{\ell(R/\mathfrak{a}_{m})}{m^{n}/n!}=\lim_{\mathbb{Z}_{>0}\ni m\to\infty}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}},$
@@
(1.1) $d_{1}(\mathfrak{a}_{\bullet},\mathfrak{b}_{\bullet})\coloneqq 2\mathrm{e}(\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet})-\mathrm{e}(\mathfrak{a}_{\bullet})-\mathrm{e}(\mathfrak{b}_{\bullet}),$
where $\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}$ is defined termwise. (1.1) is a similar expression to the Darvas metrics in the global case [Dar15, BJ25].
An issue is that we can have $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{b}_{\bullet})$ for $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$, which means that $d_{1}$ is not a metric on $\mathrm{Fil}_{R,\mathfrak{m}}$. An observation of [BLQ24] is that this issue can be resolved by considering saturated filtrations, that is, $\mathfrak{a}_{\bullet}=\widetilde{\mathfrak{a}_{\bullet}}$, where $\widetilde{\mathfrak{a}_{\bullet}}$ is the saturation of $\mathfrak{a}_{\bullet}$ defined by
[…]
for any $\lambda\in\mathbb{R}_{>0}$. Denote by $\mathrm{Fil}^{s}_{R,\mathfrak{m}}\subset\mathrm{Fil}_{R,\mathfrak{m}}$ the subset consisting of saturated filtrations. Note that there is a canonical element $\widetilde{\mathfrak{m}^{\bullet}}\in\mathrm{Fil}^{s}_{R,\mathfrak{m}}$, where $\mathfrak{m}^{\bullet}\coloneqq\{\mathfrak{m}^{\lfloor\lambda\rfloor}\}_{\lambda\in\mathbb{R}_{>0}}$ is the $\mathfrak{m}$-adic filtration.
Given $\mathfrak{a}_{\bullet,0},\mathfrak{a}_{\bullet,1}\in\mathrm{Fil}_{R,\mathfrak{m}}$, the geodesic $\mathfrak{a}_{\bullet,t}$ between them, introduced in [XZ21, BLQ24], is a segment of $\mathfrak{m}$-filtrations for $t\in[0,1]$, defined by
(1.2) $\mathfrak{a}_{\lambda,t}\coloneqq\sum_{\mu,\nu\geq 0,\ (1-t)\mu+t\nu=\lambda}\mathfrak{a}_{\mu,0}\cap\mathfrak{a}_{\nu,1}$
for any $\lambda>0$. The definition is a local analogue of the geodesics studied in [BLXZ21].
Our first main result is that the Hausdorff quotient of $(\mathrm{Fil}_{R,\mathfrak{m}},d_{1})$ is naturally $(\mathrm{Fil}^{s}_{R,\mathfrak{m}})$, which is endowed with the structure of a geodesic metric space by the above constructions.
@@
Let $(R,\mathfrak{m})$ be a Noetherian local domain that is analytically irreducible. Denote the set of all linearly bounded $\mathfrak{m}$-filtrations on $R$ by $\mathrm{Fil}_{R,\mathfrak{m}}$ and let $\mathrm{Fil}^{s}_{R,\mathfrak{m}}\subset\mathrm{Fil}_{R,\mathfrak{m}}$ be the subset consisting of saturated ones. Then
@@
    the function $d_{1}$ given by (1.1) is a pseudometric on $\mathrm{Fil}_{R,\mathfrak{m}}$,
@@
For more details about saturated filtrations, see section $2.$ For definitions related to metric spaces, see section 2.6.
@@ 1.1.2. The toric case
Let $\mathbbm{k}$ be an algebraically closed field of characteristic 0. Let $N$ be a free abelian group of rank $n\geq 1$ and $M=N^{*}$ its dual. Let $\sigma\subset N_{\mathbb{R}}=N\otimes_{\mathbb{Z}}\mathbb{R}$ be a strongly convex rational polyhedral cone of maximal dimension. We have an affine normal toric variety $X_{\sigma}=\mathrm{Spec}R_{\sigma}=\mathrm{Spec}\mathbbm{k}[\sigma^{\vee}\cap M]$ with a unique torus invariant point $x$, where $\sigma^{\vee}\subset M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$ is the dual cone of $\sigma$. Let $R$ be the local ring of $X$ at $x$ and $\mathfrak{m}$ its maximal ideal.
Let $\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}}\subset\mathrm{Fil}^{s}_{R,\mathfrak{m}}$ be the subset of all monomial filtrations (i.e. filtrations $\mathfrak{a}_{\bullet}$ such that every $\mathfrak{a}_{\lambda}$ is a monomial ideal). Via the Okounkov body construction, we identify the metric space $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ with a subspace of the space of cobounded subsets of $\sigma^{\vee}$ with the symmetric difference metric (sometimes also called the Fréchet-Nykodym-Aronszajn metric).
@@
    the metric space $(\mathrm{Fil}^{s,\mathrm{mon}}_{R,\mathfrak{m}},d_{1})$ is complete.
In the above theorem, $\mathcal{P}(\sigma^{\vee})$ denotes the set of all closed convex subsets of $\sigma^{\vee}$ with bounded complement, and the metric $d$ is defined to be the Euclidean volume of the symmetric difference of two subsets. For more details, we refer to Section 3.4. Note that $P\mapsto\chi_{P}$ gives an embedding $\mathcal{P}(\sigma^{\vee})\to L^{1}_{\mathrm{loc}}(\sigma^{\vee})$, where $\chi_{P}$ is the characteristic function of $P\subset\sigma$.
@@ 1.1.3. The supnorm metric and homogeneous norms
As in the global case [BJ25], there is also a correspondence between filtrations and norms in the local setting. Indeed, given $\mathfrak{a}\in\mathrm{Fil}_{R,\mathfrak{m}}$, we can define a function $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ on $R$ by $\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f)\coloneqq\sup\{\lambda\in\mathbb{R}\mid f\in\mathfrak{a}_{\lambda}\}$. In particular, the canonical saturated filtration $\widetilde{\mathfrak{m}^{\bullet}}$ defines a canonical norm $\chi_{0}\coloneqq\mathrm{ord}_{\widetilde{\mathfrak{m}^{\bullet}}}$. See Definition 2.7 for the definition of norms and Definition-Lemma 2.8, Lemma 2.12 and Lemma 2.29 for the detailed dictionary.
Despite the equivalence, it is sometimes more convenient to use one of the languages. As an example, we now consider another pseudometric $d_{\infty}$ on the space $\mathcal{N}_{R,\mathfrak{m}}$ of $\mathfrak{m}$-norms on $R$, which behaves like the uniform metric on the space of continuous functions on a compact space. This can also be viewed as the local analogue of the $d_{\infty}$ metric introduced in [BJ25].
@@
(1.3) $d_{\infty,\rho}(\chi,\chi^{\prime})\coloneqq\limsup_{\lambda\to\infty}\sup_{\rho(f)\geq\lambda}\frac{|\chi(f)-\chi^{\prime}(f)|}{\rho(f)}.$
@@
1.2. Weak topologies and semi-continuity of log canonical thresholds
@@
Another point of view is to consider a filtration $\mathfrak{a}_{\bullet}$ as a function on certain valuation spaces by $v\mapsto v(\mathfrak{a}_{\bullet})$, where $v(\mathfrak{a}_{\bullet})\coloneqq\lim_{\lambda\to\infty}\frac{v(\mathfrak{a}_{\lambda})}{\lambda}$. We call the topology defined this way using the space $\mathrm{Val}^{+}_{R,\mathfrak{m}}$ of valuations with positive volume the ++-topology.
We prove that the log canonical threshold satisfies certain semi-continuity on the space of filtrations, and that it is locally Lipschitz continuous with respect to the $d_{\infty}$-topology, which is similar to [DK01, Theorem 3.3].
@@
Let $(X,x)=(\mathrm{Spec}R,\mathfrak{m})$ be a klt singularity over a field $\mathbbm{k}$ of characteristic 0. Then
@@
    if $\mathfrak{a}_{\bullet,k}\in\mathrm{Fil}^{s}$ converges to $\mathfrak{a}_{\bullet}\in\mathrm{Fil}^{s}$ in the ++-topology, then
@@ 1.3. Miscellaneous results about filtrations
The spaces of filtrations are equipped with a natural partial order by inclusion, that is, $\mathfrak{a}_{\bullet}\subset\mathfrak{b}_{\bullet}$ if and only if $\mathfrak{a}_{\lambda}\subset\mathfrak{b}_{\lambda}$ for any $\lambda\in\mathbb{R}_{>0}$. One can similarly define the intersection of two filtrations termwise, and in Definition-Lemma 2.3 (Definition-Lemma 2.22, resp.), we define the join $\vee$ (saturated join $\vee_{s}$, resp.), which is the analogue of the sum of two ideals. The next theorem asserts that these operations define natural lattice structures on the spaces of $\mathfrak{m}$-filtrations.
@@
In the above theorem, $\widetilde{\mathfrak{a}^{\bullet}}$ denotes the saturation of the $\mathfrak{a}$-adic filtration $\mathfrak{a}^{\bullet}=\{\mathfrak{a}^{\lceil\lambda\rceil}\}$. The above theorem, together with Theorem 1.1, can be viewed as a higher-dimensional analogue of the results of [FJ04].
We will frequently use the following characterization for saturated filtrations, which generalizes results of [Ree56a, Ree56b].
@@
We remark that the above proposition can be viewed as an analogue of the Berkovich Maximum Modulus Principle. See also Lemma 2.32.
As an application, when $\mathfrak{a}$ is an $\mathfrak{m}$-primary ideal, we prove that the saturation of the $\mathfrak{a}$-adic filtration is determined by the set of Rees valuations of $\mathfrak{a}$, as expected.
@@
Organization of the paper. In Section 2, we recall some definitions and basic properties that will be needed. Along the way, we also prove some basic facts about filtrations. In Section 3, we introduce the Darvas metric $d_{1}$ and consider its properties. In Section 4, we define the metric(s) $d_{\infty}$, and compare the different topologies on the space of filtrations. In Section 5 we prove the main results Theorem 1.1 through Theorem 1.5. In Section 6, we discuss the relation of the results in this paper with results in the global case, give several further examples, and propose some related questions.
@@
Acknowledgments. A large portion of this paper is derived from the author’s doctoral thesis. The author would like to thank his advisor, Professor Chenyang Xu, for his constant support, encouragement and numerous inspiring conversations. Part of this work is inspired by an earlier collaboration with Harold Blum and Yuchen Liu. The author owes them special thanks for their insights on the topic.
The author would like to thank Zhiyuan Chen, Colin Fan, Yujie Luo, Tomasso de Fernex, Rémi Reboulet, Xiaowei Wang and Ziquan Zhuang for fruitful discussions. The author would also like to thank Tamás Darvas, Jingjun Han, Jihao Liu, Minghao Miao, Junyao Peng, Linsheng Wang, Lingyao Xie, Qingyuan Xue, Tong Zhang and Junyan Zhao for kind comments. Special thanks are also due to Rémi Reboulet and Mattias Jonsson for reviewing a preliminary version of this paper and offering valuable suggestions. Part of this work is done while the author visited Northwestern University, the University of Utah, Fudan University and BICMR at Peking University. The author would like to thank them for their hospitality and the amazing environment they provided.
We are also grateful to the referees for the corrections and many valuable comments, which improved the quality of this paper.